40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 20 00:19:23 2013, rana, Update, LSC, PRMI / DRMI investigations 6x
    Reply  Tue Aug 20 01:28:47 2013, Jenne, Update, LSC, REFL investigations 
       Reply  Tue Aug 20 11:41:30 2013, Koji, Update, LSC, REFL investigations 
          Reply  Tue Aug 20 11:52:20 2013, Jenne, Update, LSC, REFL investigations 
             Reply  Tue Aug 20 18:42:57 2013, Jenne, Update, LSC, REFL investigations 
Message ID: 9043     Entry time: Tue Aug 20 18:42:57 2013     In reply to: 9041
Author: Jenne 
Type: Update 
Category: LSC 
Subject: REFL investigations 

I have done the swap in the REFL path.  First, I swapped the positions of REFL11 and REFL55.  Then, I swapped out the 50/50 BS for a 90% reflection BS.  (90% goes to REFL55, 10% goes to REFL11).  I also changed the aluminum dump that was dumping the old REFL165 path into a razor dump.

Before: REFL11 had 4.0mW, REFL55 had 3.1mW.  Now, REFL11 has 0.53mW, and REFL55 has 6.9mW.  REFL165 still has around 61mW of light, and REFL33 has 3.3mW (the things that were changed were after 165 and 33 in the REFL path). 

Now, the DC value of the REFL PDs are:  REFL165 = 10.4V, REFL33 = 110mV, REFL55 = 232mV, REFL11 = 18.6mV. 

As I was finishing aligning the beams onto all of the REFL diodes, Manasa asked for the IFO so she and Masayuki could continue their work on the Xarm, so I'll check the signals acquired a little later.

ELOG V3.1.3-